OpenStudy (anonymous):

(3)to the third power(-3)to the third power

OpenStudy (stormfire1):

\[3^3 * 3^{-3}\]?

OpenStudy (stormfire1):

is that the question?

OpenStudy (anonymous):



OpenStudy (stormfire1):

\[3^3 = 9\]\[3^{-3}=1/3^3=1/9\] If you multiply them together, you get 1.

OpenStudy (anonymous):

wait thats not right

OpenStudy (stormfire1):

what's the right problem ? Just draw it :)

OpenStudy (stormfire1):

The one you drew already would be: \[3^3 = 9\]\[-3^3 = -27\]\[9 * -27 = -243\]

OpenStudy (anonymous):

um here are my options -400 -648 98 288 -729 -648